OpenStudy (anonymous):

For what number(s) x, with -2π< x < 2π, does sec(x +π/7) = 1?

hero (hero):

\(\sec(x +\dfrac{π}{7}) = 1\) Take inverse secant of both sides: \(x + \dfrac{\pi}{7} = \sec^{-1}(1)\) \(x = \sec^{-1}(1) - \dfrac{\pi}{7}\)

hero (hero):

There's apparently more than one solution.

OpenStudy (anonymous):

yea theres ment to be more than one solution

OpenStudy (anonymous):

could you please help, i dont knw how to work it out??

hero (hero):

Finding the rest of them has to do with the period of the secant function

hero (hero):

Do you know the period of secant function?

OpenStudy (sweetburger):

i recommend memorizing the periods of the main functions also i think you can solve for them using the unit circle but i could be wrong

hero (hero):

There's one other solution that is most obvious if you know the period of the function